OpenStudy (anonymous):

the length of a container is 2dm more than the width. the width is 6dm. the height is twice the length. how much liquid does it contain if half-filled?

OpenStudy (anonymous):

The half volume evaluated at w =6 is:\[\frac{w(w+2)(2(w+2))}{2}=w (2+w)^2=6 (2+6)^2=384 \text{dm}^3 \]